Melanie can paint a certain barn by herself in X days.
Her helper Melissa can paint the same barn by herself in 2x days.
Write a rational expression for the fraction of the barn,
that they complete in one day by working together. Evaluation the
expression for X=5.. Can you help me?

Answer by ewatrrr(24785)


Hi,
Melanie can paint a certain barn by herself in X days. Milissa in 2x days.
the fractional part of the barn,that they complete in one day by working together.
1/x + 1/2x
x = 5
1/5 + 1/10 = 2/10 + 1/10 = 3/10 of the barn can be painted in one day working together.
(3 1/10 days to finish it)